A single phase overhead transmission line delivers 4000 kW at 11 kV at 0·8 p.f. lagging. If resistance and reactance per conductor are 0·15 Ω and 0·02 Ω respectively, calculate

: (i) percentage regulation

(ii) sending end power factor

(iii) line losses 


1
Expert's answer
2021-01-20T04:56:25-0500

Is = Ir = 454.56(L-36.86)

i)- percentage regulation = "\frac{(Vs-Vr)}{Vr}\times 100"

="\frac{(11120- 110000)}{110000}\times 100"

= 1.09%

ii) "cos\phi= Cos(36.86- 0.3465)"

= 0.80log

iii) line losses = "I^2Rl"

= "(454.54)^2(.30)"

losses = 61982W